The length of a rectangle is 2 in. more than twice its width. If the
perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by checkley75(3666)

2X+2(2+2X)=34
2X+4+4X=34
6X=34-4
6X=30
X=30/6
X=5 LENGTH OF THE SHORTER SIDE.
2+2*5
2+10
12 IS THE LENGTH OF THE LONGER SIDE.
PROOF
2*5+2(12)=34
10+24=34
34=34